OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (hoblos):

y = -3x + 5 y = 2x + 5 so -3x + 5 = 2x + 5 -5x = 0 x=0 y=5 so (0, 5) D